# Moral Alignment for LLM Agents

### v3

@@ 6 Discussion
In this work, we present a method for fine-tuning LLM agents to adhere to a specific moral strategy in matrix games by employing RL with intrinsic rewards.
The two different moral payoff structures used in this study have different advantages and disadvantages in terms of implementation in real-world systems. Our definition of the consequentialist (Utilitarian) moral agents is a function of the payoffs given by the environment to both players. Thus, its implementation in practice requires that the LLM agent will has observability of the rewards received by both players from the environment on a given time step (or a reliable estimate). For Deontological morality, on the other hand, a norm may be easier to define in any environment without direct access to game payoffs or the opponent’s current actions. The definition of the Deontological norm used in this study (“do not defect against a cooperator”) only required a memory of one previous move of an opponent. This makes such a norm-base reward function easy to implement in cases in which the developer of an LLM agent only has access to their own agent’s observations of the environment and not anyone else’s. In future work, the intrinsic rewards approach can be applied to modeling a variety of other moral values.
An extension of this method to other environments would be of great interest, including fine-tuning agents with other payoff structures, more complex games or longer history lengths (for example, to aid the development of continually-learning LLM agents in practice), as well as text-based scenarios that tap into a variety of moral values. Furthermore, the method of intrinsic rewards could also be applied in a multi-objective manner to address the issue of pluralistic alignment (Sorensen et al., 2024) - in particular, a single agent could be trained with a combination of rewards representing different moral values. This may provide a promising direction for building agents that are able to satisfy the moral preferences of a wide range of individuals, which currently remains an open problem in alignment (Anwar et al., 2024; Ji et al., 2024). Finally, agents trained via intrinsic rewards as proposed in this study could also form the basis for a Constitutional AI architecture composed of artificial agents characterized by different moral frameworks (Bai et al., 2022).

### v4

@@ 6 Discussion
In this work, we present a method for fine-tuning LLM agents to adhere to a specific moral strategy in matrix games by employing RL with intrinsic rewards. The two core moral payoff structures used in this study have different advantages and disadvantages in terms of implementation in real-world systems. Our definition of the consequentialist (Utilitarian) moral agents is a function of the payoffs given by the environment to both players. Thus, its implementation in practice requires that the LLM agent has observability of the rewards received by both players from the environment on a given time step (or a reliable estimate). For Deontological morality, on the other hand, a norm may be easier to define in any environment without direct access to game payoffs or the opponent’s current actions. The definition of the Deontological norm used in this study (“do not defect against a cooperator”) only requires a memory of one previous move of an opponent. This makes such a norm-base reward function easier to implement in cases in which the LLM agent only has access to their own observations of the environment and not anyone else’s. In future work, the intrinsic rewards approach can be applied to modeling a variety of other moral values.
An extension of this method to other environments would be of great interest, including fine-tuning agents with other payoff structures, more complex games or longer history lengths (for example, to aid the development of continually-learning LLM agents in practice), as well as text-based scenarios that tap into a variety of moral values. Furthermore, the method of intrinsic rewards could also be applied in a multi-objective manner to address the issue of pluralistic alignment (Sorensen et al., 2024) - in particular, a single agent could be fine-tuned with a combination of rewards representing different moral values. This may provide a promising direction for building agents that are able to satisfy the moral preferences of a wide range of individuals, which currently remains an open problem in alignment (Anwar et al., 2024; Ji et al., 2024). Finally, agents trained via intrinsic moral rewards as proposed in this study could also form the basis for a Constitutional AI architecture composed of artificial agents characterized by different moral frameworks (Bai et al., 2022).
@@ 7 Conclusion
Acknowledgments
This work was partially supported by the Leverhulme Trust through the Doctoral Training Programme for the Ecological Study of the Brain - DS-2017-026 (Elizaveta Tennant), and partially supported by the Italian Ministry of University and Research (MUR) through the project PRIN 2022 “Machine-learning based control of complex multi-agent systems for search and rescue operations in natural disasters (MENTOR)” - CUP E53D23001160006 (Mirco Musolesi and Elizaveta Tennant).